How do you find the measurement of 2d shapes?
1
Expert's answer
2020-05-11T12:07:03-0400

Rectangle



a, b - sides.

A, B, C, D - angles

"\angle A= \angle B=\angle C = \angle D = 90 ^\degree"

area "S = a * b"

perimeter P = 2 (a + b)


Parallelogram



a, b - sides.

A, B, C, D - angles

"\angle A= \angle C" "\angle B = \angle D"

area "S = b * h"

perimeter "P = 2 (a + b)"


Trapezoid


a, b, c, d - sides

A, B, C, D - angles

area "S= \frac{b+c}{2}h"

perimeter "P = a + b + c + d"


Triangle



a, b, c - sides

A, B, C- angles

area "S=\frac{1}{2}ch"

perimeter "P=a+b+c"


Ellipse




area "S=\pi ab"

perimeter "P=\pi \frac{a+b}{2}"